Write an equation in point-slope form of the line with slope –8 that contains the point J(–3, 2).
Answer by ewatrrr(24785)


Hi,
equation in point-slope form of the line with slope –8 containing J(-3,2)
(y - 2) = -8(x + 3)
tranlates to: y = -8x -22 in the slope-intercept form
the standard slope-intercept form for an equation of a line is y = mx + b
where m is the slope and b the y-intercept.